The number of k-Dimensional Partitions for Seven Through Twelve

Shalosh B. EKHAD ^{1}

References

[A] G. E. Andrews, “The Theory of Partitions”, Addison-Wesley, 1976. Reprinted by Cambridge University Press, 1984. First paperback edition, 1998.

[ABMM] A.O.L. Atkin, P. Bratley, I.G. Macdonald, and J.K.S. McKay, Some Computations for m-dimensional partitions, Proc. Cambridge Phil. Soc. 63 (1967), 1097-1100.

[BGP] S. Balakrishnan, S. Govindarajan and N. S. Prabhakar, On the asymptotics of higherdimensional partitions, arXiv:1105.6231 .

1 ShaloshBEkhad@gmail.com .

Exclusively published in: http://www.math.rutgers.edu/~zeilberg/pj.html and http://arxiv.org/ .

Feb. 15, 2012.